OpenStudy (anonymous):

A type of weeping cherry tree will flower when it is taller than140 cm but no more than 200 cm. If a tree is purchased when it is 50 cm tall and it grows 20 cm each year, during what time period will the tree flower? List the lower and upper limits seperated by a comma. ex: (1,5)

OpenStudy (anonymous):

Height of tree = Start height + Heightrate * years H = 50 + 20x (if x is the amount of years) if H = 140: 140 = 50 + 20x, x = 4.5 years if H = 200 cm: 200 = 50 + 20x, x = 7.5 years (4.5, 7.5)